OpenStudy (anonymous):

3.05]Which is less than zero point six five? thirteen-twentieths fourteen twenty-fifths four-fifths seven-tenths

OpenStudy (anonymous):

I would find the fractional form of 0.65 which is 65/100 and now get them all under a common denominator, 100 in this case and compare the numerators

OpenStudy (mathstudent55):

Compare 0.65 wirth each number below. 13/20 = 0.65 14/25 = 0.56 4/5 = 0.8 7/10 = 0.7

OpenStudy (anonymous):

so its seven - tenths

OpenStudy (anonymous):

so wats the answer for my sis trying to prove her wrong

OpenStudy (mathstudent55):

You need an answer that is _less- than 0.65. 0.65, 0.56, 0.80, 0.70 Which of the 4 above is _less_ than 0.65?

OpenStudy (anonymous):

four - fifths

OpenStudy (anonymous):

can u tell me the answer my sis is getting antous

OpenStudy (mathstudent55):

All numbers are in hundredths. Can you tell which number is less than 65? 65, 56, 80, 70?

OpenStudy (anonymous):

56

OpenStudy (mathstudent55):

Right, so since all numbers you have are in hundredths, 0.56 is the only one that is less than 0.65.

OpenStudy (mathstudent55):

And 0.56 is 14/25

OpenStudy (anonymous):

thx I have another question I need help on for my sister

OpenStudy (mathstudent55):

Ok

OpenStudy (anonymous):

3.03]Which can not be written as an equivalent improper fraction? two-thirds twenty-two fourths eleven-sixths four and one-eighth

OpenStudy (mathstudent55):

An improper fraction is a fraction whose numerator is greater than its denominator. In a fraction, the number on top is the numerator, and the number on the bottom is the denominator. Any mixed numeral can be written as an improper fraction.

OpenStudy (anonymous):

im cofused then @mathstudent55

OpenStudy (mathstudent55):

Look at your choices: 

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

is the answer

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

flvs.net